A total of $66,000 is to be​ invested, some in bonds and some in certificates of deposit​ (CDs). If the amount invested in bonds is to exceed that in CDs by
$1,000, how much will be invested in each type of​ investment

Answer by rfer(16322)

x+x+1000=66000
2x=65000
x=65000/2
x=32500
x+1000=33500